OpenStudy (anonymous):

3/3sqrt2-sqrt3 rationalize the denominator and simplify can i get an explanation please

OpenStudy (anonymous):

Is your original expression: a) \(\frac{3}{3\sqrt{2} - \sqrt{3}}\) b) \(\frac{3}{3\sqrt{2}} - \sqrt{3}\)

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Ok, so multiply top and bottom by the conjugate \(3\sqrt{2} + \sqrt{3}\)

OpenStudy (anonymous):

top and bottom??

OpenStudy (anonymous):

yes, anything you do to the denominator you must do to the numerator to keep the ratio the same.

OpenStudy (anonymous):

Anything you multiply the denominator by that is. \[\frac{4}{3} = \frac{4*2}{3*2} = \frac{8}{6}\]

OpenStudy (anonymous):

im still confused i wrote this \[3/3\sqrt2-\sqrt3 *3\sqrt2+\sqrt3/3\sqrt2+\sqrt3\]

OpenStudy (anonymous):

\[3/3\sqrt2-\sqrt3 * 3\sqrt2+\sqrt3/3\sqrt2+\sqrt3\]i mean

OpenStudy (anonymous):

You need to use parentheses

OpenStudy (anonymous):

\[\frac{3}{3\sqrt{2} - \sqrt{3}} = \frac{3(3\sqrt{2} + \sqrt{3})}{(3\sqrt{2} - \sqrt{3})(3\sqrt{2} + \sqrt{3})}\]

OpenStudy (anonymous):

And on bottom we get some nice simplification.

OpenStudy (anonymous):

i dont understand how that simplifies

OpenStudy (anonymous):

Foil it out.. \[(3\sqrt{2} - \sqrt{3})(3\sqrt{2} + \sqrt{3})\] \[=3\sqrt{2}*3\sqrt{2} - 3\sqrt{2}\sqrt{3} +3\sqrt{2}\sqrt{3} -\sqrt{3}*\sqrt{3}\] \[=9(2) + 0 - 3 = 15\]

OpenStudy (anonymous):

Which cancels with the 3 up top

OpenStudy (anonymous):

to give you \[\frac{3\sqrt{2} + \sqrt{3}}{5}\]

OpenStudy (anonymous):

i am not grasping this

OpenStudy (anonymous):

i dont understand how you got 15

OpenStudy (anonymous):

i dont understand how you got 15

OpenStudy (anonymous):

This is what you had in the denominator right? \[(3\sqrt{2} - \sqrt{3})(3\sqrt{2} + \sqrt{3})\] This is just like \[(a - b)(a+b) = a^2 -ba + ab - b^2\] The middle two terms cancel each other out and you are left with \[a^2 - b^2\] Where in your case, \(a=3\sqrt{2}\) and \(b=\sqrt{3}\) So: \[a^2 - b^2 \]\[= (3\sqrt{2})^2 - (\sqrt{3})^2 \]\[= 3^2\sqrt{2}^2 - \sqrt{3}^2 \]\[=9(2) - 3 = 18-3 = 15\]

OpenStudy (anonymous):

i understand the foil method better thankyou i understand now